-4x-2y=-36
4x+y=26

Answer by checkley79(3341)

-4x-2y=-36
4x+y=26 ADD THESE 2 EQUATIONS.
----------
-Y=-10
Y=10 ANS.
-4X-2*10=-36
-4X-20=-36
-4X=-36+20
-4X=-16
X=-16/-4
X=4 ANS.
PROOF:
4*4+10=26
16+10=26
26=26